What is the amount of o2 liberated if 9g of water is taken with 50%current efficiency ?
 

Grade:12th pass

1 Answers

Arun
5 years ago
H2O.................> H2..+(1/2) * O29g1/2 mole.........1/2mol.......1/4 molefor ¼ mole of O2, volume=22.4*1/4=5.6 L at STP